How do we compare correlation to regression?



How do we calculate regression?


1
Expert's answer
2021-11-16T11:07:36-0500

Correlation is a statistical measure that determines the association or co-relationship between two variables. Regression describes how to numerically relate an independent variable to the dependent variable.


regression equation:

"y=ax+b"

where


"a=\frac{n\sum xy-\sum x\sum y}{n\sum x^2-(\sum x)^2}"


"b=\frac{\sum y\sum x^2-\sum x\sum xy}{n\sum x^2-(\sum x)^2}"